# Characteristic Root Analysis and Regularization for Linear Time Series Forecasting

### v2

@@ E.1 Dataset
For the synthetic dataset we used in Section 5.3 to study data scaling property and noise robustness, the train, validation, and test sets are split with a $5:2.5:2.5$ ratio.
@@
Additionally, we also use M4 [Makridakis et al., 2020] dataset to test the robustness of our methods under short lookback window as well as their performances in short-term forecasting tasks. The M4 dataset consists of 100,000 real-world time series from diverse domains such as finance, economics, demographics, and industry. Each subset of M4 is collected in yearly, quarterly, monthly, weekly, daily, and hourly frequencies, with a mixture of data from the domains mentioned before. For the M4 dataset, the lookback window $L$ is set to $2\times H$, where $H$ is the forecasting horizon, which differs from long-term forecasting datasets. The details of M4 dataset are listed in Tabel 7.
@@ E.2 Implementation Details
For more details on the running settings of Root Purge, please refer to our code and running scripts in the supplementary material  [^9] Upon acceptance, we will open source our code..
@@ E.12 Additional Results on Data Scaling & Noise Robustness on Simulation Data
E.13 Root Analysis on Simulation Data
In most real-world applications, obtaining noise-free system dynamics is challenging, if not impossible. Therefore, following the principle that accurately identifying the characteristic roots is sufficient for minimizing forecasting error (as established in Section 3, Fact 1), we have used Mean Squared Error (MSE) as the primary benchmark for our methods. In this section, however, we directly evaluate how Rank Reduction Regression (RRR) and Root Purge affect the learned root distribution in a synthetic setting. This allows us to validate that our methods indeed recover roots that are closer to the true, noise-free roots.
Such a direct comparison is infeasible with real-world data due to the lack of access to the underlying noise-free dynamics and their true characteristic roots. To address this, we conduct experiments on synthetic data generated by the following dynamics:
@@ E.13 Root Analysis on Simulation Data
We use the same dynamics from Section 5.3, which include periodic components and a linear trend; this specific form is not critical, as similar results are obtained with other dynamics. The term $\sigma\cdot\varepsilon_{t}$ represents additive Gaussian noise. In the noise-free case ($\sigma=0$), this signal can be exactly represented using six characteristic roots. For clearer visualization, we fit a linear model with a lookback window of length 25 and a forecasting horizon of length 25 to limit the number of roots. We compare the 25 characteristic roots learned under four conditions against the ground truth roots:
@@
    A standard linear model is fitted using Ordinary Least Squares (OLS) on the noisy signal, and the root distribution from the first forecasting head is visualized.
@@
    Our Root Purge method is applied to the same noisy data, and the resulting root distribution is plotted.
@@
    Similarly, our Rank Reduction Regression (RRR) approach is applied, and the learned roots are examined.
@@
Figure 18: Visualization of root distribution on the (a) 11st and (b) $12$th linear time series forecasting horizons, as defined by Section 3, Equation (2).Both RRR and Root Purge yield root distributions visibly closer to the noise-free ground truth than the standard linear model.
We visualize the characteristic roots from these models in the two-dimensional complex plane in Figure 18. The roots obtained via RRR and Root Purge are visibly closer to the ground truth than those from the standard model. Furthermore, we quantitatively evaluate the root recovery performance over five experimental runs. For each model, we compute the mean and standard deviation of the distances between all learned roots and the ground truth roots, with root pairs matched using the Hungarian algorithm [Kuhn, 1955]. The results are summarized in Table 20.
@@
These results demonstrate that our proposed techniques enhance the robustness of root learning in the presence of noise. They provide direct empirical evidence supporting our claim that RRR and Root Purge help the model avoid capturing spurious roots, thereby strengthening the link between our theoretical insights and the observed learning behavior.

### v3

@@ E.1 Dataset
For the synthetic dataset we used in Section 5.3 to study data scaling property, noise robustness, and root distribution, the train, validation, and test sets are split with a $5:2.5:2.5$ ratio.
@@
Additionally, we also use M4 [Makridakis et al., 2020] dataset to test the robustness of our methods under short lookback window as well as their performances in short-term forecasting tasks. The M4 dataset consists of 100,000 real-world time series from diverse domains such as finance, economics, demographics, and industry. Each subset of M4 is collected in yearly, quarterly, monthly, weekly, daily, and hourly frequencies, with a mixture of data from the domains mentioned before. For the M4 dataset, the lookback window $L$ is set to $2\times H$, where $H$ is the forecasting horizon, which differs from long-term forecasting datasets. The details of M4 dataset are listed in Tabel 8.
@@ E.2 Implementation Details
For more details on the running settings of Root Purge, please refer to our code and running scripts in the supplementary material [^10] Our code is available at https://github.com/Wangzzzzzzzz/Root-Purge-for-Time-Series-Forecasting.
@@ E.12 Additional Results on Data Scaling & Noise Robustness on Simulation Data
E.13 Details of Root Analysis on Simulation Data and Quantitative Evaluation
As mentioned in the main text, a direct root comparison is infeasible with real-world data due to the lack of access to the underlying noise-free dynamics and their true characteristic roots. We employ the same synthetic dynamics used previously to illustrate the data scaling and noise robustness properties of our methods.
@@ E.13 Details of Root Analysis on Simulation Data and Quantitative Evaluation
The additive Gaussian noise term is given by $\sigma\cdot\varepsilon_{t}$. In the noise-free case ($\sigma=0$), the signal can be represented exactly using six characteristic roots. We fit a linear model with a lookback window and a forecasting horizon each of length 25—this limits the total number of learned roots and offers sufficient capacity to capture the noise-free dynamics exactly. We then compare the characteristic roots estimated under four conditions against the ground truth roots.
@@
    A standard linear model is fitted using Ordinary Least Squares (OLS) on the noisy signal, and the root distribution from the first and the twelfth forecasting head is studied.
@@
    Our Root Purge method is applied to the same noisy data, and the learned roots are examined.
@@
    Similarly, our Rank Reduction Regression (RRR) approach is applied. The learned roots are examined.
@@
Figure 18: Root distribution on the (a) 11st and (b) $12$th linear time series forecasting horizons, as defined by Section 3, Equation (2). Both RRR and Root Purge yield root distributions visibly closer to the noise-free ground truth than the standard linear model. The arrows denote the ground truth roots toward which the respective methods should converge.
In section 5.3, we see that roots obtained via RRR and Root Purge are closer to the ground truth than those from the standard model. Here, we qualitatively show the root recovery performance in Figure 18 by visualizing the characteristic roots from these models in the two-dimensional complex plane. The roots obtained via RRR and Root Purge are visibly closer to the ground truth than those from the standard model.
Additionally, since Root Purge is optimized with gradient method, we also introduced “Standard Linear Model (gradient method)" as a further baseline in Table 21 to investigate the impact of using gradient based-optimizations.
@@
These results align with the results in Section 5.3 and demonstrate that our proposed techniques enhance the robustness of root learning in the presence of noise.
E.14 Stop-Gradient and Full-Gradient Root Purge
In Appendix D.7, we analyzed the properties of the two variants of Root Purge from a theoretical perspective. Our results indicate that their theoretical behavior should be highly similar. In this section, we empirically compare the full-gradient version with the stop-gradient version. The results, summarized in Table 22 for the ETT datasets, show that both variants yield very similar performance. Therefore, we primarily use the stop-gradient version in our experiments, as it greatly simplifies gradient computation and is significantly less memory-intensive than the full-gradient version.
Table 22: Experimental results comparing the stop-gradient and full-gradient versions of Root Purge on ETT datasets with different $\lambda$ values. The stop-gradient variant produces nearly identical performance while being more efficient.
Dataset Horizon Stop Gradient Version Full Gradient Version
ETTh1 96 $0.368\pm 0.000$ $0.364\pm 0.000$ $0.359\pm 0.000$ $0.368\pm 0.000$ $0.364\pm 0.000$ $0.359\pm 0.000$
192 $0.403\pm 0.000$ $0.399\pm 0.000$ $0.394\pm 0.000$ $0.403\pm 0.000$ $0.399\pm 0.000$ $0.394\pm 0.000$
336 $0.427\pm 0.000$ $0.425\pm 0.000$ $0.423\pm 0.000$ $0.427\pm 0.000$ $0.425\pm 0.001$ $0.423\pm 0.000$
720 $0.423\pm 0.000$ $0.421\pm 0.000$ $0.423\pm 0.000$ $0.423\pm 0.000$ $0.421\pm 0.000$ $0.421\pm 0.001$
ETTh2 96 $0.270\pm 0.000$ $0.269\pm 0.000$ $0.268\pm 0.000$ $0.270\pm 0.000$ $0.269\pm 0.000$ $0.268\pm 0.000$
192 $0.330\pm 0.000$ $0.329\pm 0.000$ $28\pm 0.000 0.330$$30\pm 0.000$ $0.329\pm 0.000$ $0.328\pm 0.000$
336 $0.354\pm 0.000$ $0.355\pm 0.000$ $0.356\pm 0.000$ $0.354\pm 0.000$ $0.355\pm 0.000$ $0.356\pm 0.001$
720 $0.377\pm 0.000$ $0.377\pm 0.000$ $0.379\pm 0.000$ $0.377\pm 0.000$ $0.377\pm 0.000$ $0.378\pm 0.000$
ETTm1 96 $0.306\pm 0.000$ $0.306\pm 0.000$ $0.305\pm 0.000$ $0.306\pm 0.001$ $0.305\pm 0.001$ $0.305\pm 0.001$
192 $0.335\pm 0.000$ $0.334\pm 0.000$ $0.333\pm 0.000$ $0.335\pm 0.001$ $0.334\pm 0.000$ $0.333\pm 0.000$
336 $0.363\pm 0.000$ $0.361\pm 0.000$ $0.360\pm 0.000$ $0.362\pm 0.001$ $0.361\pm 0.000$ $0.360\pm 0.000$
720 $0.413\pm 0.000$ $0.412\pm 0.000$ $0.413\pm 0.000$ $0.413\pm 0.000$ $0.412\pm 0.000$ $0.412\pm 0.000$
ETTm2 96 $0.162\pm 0.000$ $0.161\pm 0.000$ $0.161\pm 0.000$ $0.161\pm 0.000$ $0.161\pm 0.000$ $0.161\pm 0.001$
192 $0.217\pm 0.000$ $0.217\pm 0.000$ $0.216\pm 0.000$ $0.216\pm 0.000$ $0.216\pm 0.000$ $0.216\pm 0.000$
336 $0.269\pm 0.000$ $0.269\pm 0.000$ $0.269\pm 0.000$ $0.269\pm 0.000$ $0.269\pm 0.001$ $0.269\pm 0.000$
720 $0.350\pm 0.000$ $0.350\pm 0.000$ $0.355\pm 0.000$ $0.350\pm 0.000$ $0.351\pm 0.000$ $0.354\pm 0.000$